### v2

# Evaluating Human-AI Collaboration: A Review and Methodological Framework

Abstract
The use of artificial intelligence (AI) in collaborative working environments, termed Human-AI Collaboration (HAIC), has become vital across diverse domains, enhancing decision-making, efficiency, creativity, and innovation. Despite its wide potential, evaluating HAIC effectiveness poses significant challenges due to the complex, dynamic, and reciprocal interactions between humans and AI systems. This paper provides a comprehensive analysis of existing HAIC evaluation methods and introduces a novel, domain-agnostic framework for systematically assessing these systems. To support structured assessment, our framework incorporates a decision-tree-based approach that helps select relevant metrics tailored to distinct HAIC modes (AI-Centric, Human-Centric, and Symbiotic). By integrating both quantitative and qualitative evaluation criteria, the framework aims to address key gaps in HAIC assessment methodologies. We highlight the unique challenges posed by evaluating creative and linguistic AI applications, such as large language models and generative AI in the arts, underscoring the need for tailored evaluation approaches in these emerging areas. This work lays the foundation for future research and empirical validation, offering a structured methodology to enhance the evaluation of HAIC systems.
Keywords Collaboration  $\cdot$ Evaluation  $\cdot$ Symbiosis  $\cdot$ Metrics  $\cdot$ Artificial Intelligence  $\cdot$ Interaction  $\cdot$ Adaptability
1 Introduction
Evaluating Human-AI Collaboration (HAIC) remains a significant challenge due to the complex interactions between humans and AI systems. The effectiveness of HAIC depends not only on system performance but also on the quality of human-AI interaction, trust, and adaptability. However, existing evaluation methods often fail to capture these multidimensional aspects, making it difficult to establish comprehensive assessment criteria.
To address this issue, a structured framework that provides a foundation for systematically assessing HAIC across various domains, is introduced in the paper. Is is inspired by existing HAIC evaluation methodologies, while it aims at integrating and complement them. While the framework has not yet been applied in real-world scenarios, it is designed to be adaptable and to serve as a reference for future empirical validation.
Section 2 presents a literature review on HAIC evaluation methods, summarizing different approaches, key metrics, and domain-specific challenges. This review highlights gaps in existing evaluation strategies, particularly their inability to fully capture the dynamic and reciprocal nature of HAIC interactions. Based on the finding of this review and the shortcoming identified we introduce an structured evaluation framework [bolman2013reframing] that systematically assesses HAIC across multiple domains (Sections 3– 4) . This framework integrates key evaluation factors such as goals, interaction quality, and task allocation, ensuring a standardized and adaptable assessment tool for practitioners and researchers.
1.1 Background
The rapid advancement of Artificial Intelligence (AI) has transformed multiple domains, leading to a shift from traditional Human-Machine Interaction (HMI) toward Human-AI Collaboration (HAIC) [1]. Unlike conventional HMI, which emphasizes usability and task performance [2, 3], HAIC involves a symbiotic partnership where humans and AI work together to achieve shared goals.
This paradigm shift has demonstrated immense potential across diverse sectors, including manufacturing [4, 5], financial services [6, 7], healthcare [8, 9], and education [10]. However, despite its widespread adoption, evaluating HAIC remains an unresolved challenge. Traditional assessment frameworks, which focus on efficiency and accuracy, are insufficient to capture the complexity of human-AI interaction, adaptability, and trust.
This limitation motivates the need for a comprehensive review of HAIC evaluation methodologies (Section 2) and the development of a structured framework that accounts for these multidimensional aspects (Sections 3– 4).
1.2 Challenges
While HAIC offers numerous opportunities, evaluating its effectiveness remains a complex and unresolved challenge. Recent work by Woelfle et al. (2024) [11] highlights the limitations of traditional evaluation metrics such as task performance, response time, and user satisfaction. Their study emphasizes the need for more comprehensive benchmarks that account for the dynamic and reciprocal nature of HAIC interactions, incorporating metrics like evidence appraisal quality, trust, and adaptability. Such advancements aim to better reflect the multifaceted and collaborative dimensions of human-AI partnerships.
One of the primary challenges lies in assessing the individual and joint contributions of humans and AI systems to collaborative outcomes. Unlike traditional human-machine paradigms, HAIC involves a deeper integration of human intuition and AI capabilities. Evaluating this synergy requires novel metrics that effectively capture the mutual influences between collaborators.
The unique challenges posed by creative and linguistic domains further highlight the limitations of traditional evaluation methods. Applications such as Large Language Models (LLMs) and generative AI in the arts demonstrate the transformative potential of AI to enhance creativity and human expression. However, evaluating these systems requires specialized metrics that move beyond task-based measures. For instance, assessing the interplay of human and AI contributions in artistic creation or the ethical implications of AI-generated content demands methodologies that capture qualitative and contextual nuances. While this study provides a foundational framework for HAIC evaluation, exploring the specific requirements of creative and linguistic domains is an essential avenue for future research.
Additionally, measuring the quality of human-AI interactions remains a critical yet underexplored area. Interaction quality encompasses shared decision-making processes, communication dynamics, and mutual adaptability. Evaluating the ability of both the AI system and human participants to respond to evolving circumstances is crucial for understanding the resilience and effectiveness of HAIC systems in dynamic and high-stakes environments.
1.3 Proposed Approach
Drawing inspiration from established frameworks [12, 2], our approach emphasizes ethical considerations, adaptability, and domain-specific customization. The proposed framework provides a comprehensive structure for assessing HAIC across various modes of interaction, from AI-centric to symbiotic collaborations, integrating quantitative and qualitative measures. Furthermore, it aims to assist researchers and practitioners in identifying existing gaps and opportunities for improving collaboration between humans and AI. By incorporating domain-specific insights, the framework is designed to address practical challenges and applications across fields such as healthcare, finance, and education, thereby demonstrating its versatility and relevance in diverse contexts.
1.4 Contributions
This paper makes the following key contributions:
  • A review of existing HAIC evaluation methodologies, identifying key gaps and challenges that limit their applicability across diverse domains.
  • The introducion a structured evaluation framework for HAIC, integrating both quantitative and qualitative metrics to assess performance, interaction quality, and task allocation.
  • The framework is domain-agnostic, designed to support AI-centric, human-centric, and symbiotic collaboration models, offering a flexible structure applicable to various application areas.
2 Literature Review
This section first examines the transition from HMI to HAIC, highlighting key evaluation challenges. It then explores the fundamental elements that define HAIC interactions, providing the foundation for developing a structured evaluation framework.
2.1 From Human-Machine Interaction to Human-AI Collaboration
The rapid advancement of HAIC necessitates a broader evaluation perspective that goes beyond traditional HMI usability assessments [13, 14, 15]. HMI methods have primarily been designed to measure user experience and task performance, relying on metrics such as response time, accuracy, and efficiency. However, HAIC introduces new challenges that extend beyond interface usability, requiring metrics that capture interaction quality, collaboration effectiveness, and trust.
Okamura and Yamada (2020) [16] emphasize the importance of system transparency in maintaining trust calibration between humans and AI agents. Over-trusting AI can lead to automation bias, while insufficient trust may result in underutilization. Similarly, Yue et al. (2023) [17] examine how different human-AI collaboration models influence user perceptions and adoption, reinforcing the need for evaluation methods that incorporate trust, interpretability, and adaptability.
2.2 Core Elements and Modes of HAIC
Human-AI Collaboration (HAIC) is defined by the cooperative partnership between individuals and AI systems to accomplish shared objectives. Understanding its key elements is essential for designing effective evaluation methodologies.
  • Tasks: HAIC systems support diverse tasks, ranging from automated decision-making to creative problem-solving. The nature of a given task dictates the required level of collaboration between humans and AI [18, 19]. For instance, in healthcare, an AI system may analyze medical images, while a radiologist interprets and integrates results into a final diagnosis.
  • Goals: Collaboration in HAIC is driven by individual and collective goals. AI objectives often include enhancing efficiency and accuracy, while human goals may involve skill augmentation and contextual reasoning [20]. Continuing with the radiology example, the AI assists in accurate disease detection, while radiologists focus on clinical decision-making and patient-centered care.
  • Interaction: The effectiveness of HAIC is determined by communication quality and feedback mechanisms [21, 22]. Misalignment in expectations or miscommunication can lead to ineffective collaboration. For instance, AI-driven customer support chatbots must interpret user queries and refine responses dynamically to ensure effective problem resolution.
  • Task Allocation: Dynamic task allocation is essential in HAIC, ensuring that responsibilities shift based on real-time conditions and expertise distribution [23]. For example, in advanced driver assistance systems, AI controls navigation under normal conditions, but the human driver intervenes when unexpected obstacles arise.
Figure 1 visually represents the key elements of HAIC, which are important for developing effective evaluation frameworks. By clearly defining the nature of the collaboration, the intended goals, and the dynamics of interaction, we can then create metrics and assessments that accurately measure the success and impact of HAIC systems.
Figure 1: Core elements of Human-AI Collaboration (HAIC) and examples of their application.
One can use any of the HAIC elements, i.e., tasks, goals, types of interaction, and task allocation to differentiate between different HAIC methods. For the purposes of evaluation, however, perhaps the most useful distinction is in terms of allocation [24]. We identify three modes, namely AI-centric, Human-centric, and Symbiotic, which are further described below.
Human-Centric Mode
In the Human-centric mode, humans retain the primary decision-making authority, utilizing AI as an augmentative tool to enhance human capabilities without superseding the human role. This mode values human intuition and oversight, with artificial intelligence employed to manage repetitive or data-intensive tasks. It optimizes the unique cognitive abilities of humans for strategic and ethical decision-making while leveraging AI to augment human capabilities.
Supporting this mode, Sutthithatip [25] highlights the use of Explainable AI (XAI) in high-stakes sectors like aerospace to demystify complex AI operations, thereby enhancing human decision-making. Additionally, How et al. [26] describe the integration of AI in sustainable development, which facilitates the involvement of non-experts in data-driven policy making, illustrating the supportive role of AI in enhancing human capabilities.
Symbiotic Mode
The Symbiotic mode represents a balanced partnership where humans and AI systems collaborate closely, mutually enhancing each other’s capabilities. This mode is characterized by a two-way interaction, shared decision-making, and a continuous exchange of feedback, aiming to achieve collective goals through a synergistic relationship. This mode is particularly relevant in complex tasks where both human intuition and AI’s computational capabilities are essential for optimal outcomes.
Studies like those by Cesta et al. [27] explore the symbiotic integration of AI in operational settings, enhancing the collaborative dynamics between humans and robots. Moreover, Jokinen & Watanabe [28] emphasize the crucial role of trust in these interactions, which is fundamental for a successful symbiotic relationship. Mahmud et al. [29] and Sowa et al. [30] further illustrate how symbiotic HAIC can revolutionize creative fields and managerial roles by combining human creativity with AI’s analytical strengths to foster innovation and improve decision-making.
A key example of the symbiotic mode is the Learning to Defer (L2D) paradigm, where AI systems are designed to recognize their own limitations and defer decisions to human judgment when necessary. This approach enhances collaborative decision-making by leveraging the strengths of both humans and AI. Recent studies have expanded on this concept. Alves et al. (2024) [31] introduced a cost-sensitive L2D framework that considers scenarios with multiple experts and workload constraints, emphasizing the need for flexibility in real-world Human-AI Collaboration (HAIC) systems. Additionally, Hemmer et al. (2022) [32] proposed an approach that trains a classification model to complement the capabilities of multiple human experts. Their work underscores the adaptability of the HAIC systems’ participants, showcasing the ability to dynamically modify task allocation according to the availability and expertise of various agents.
AI-centric Mode
The AI-centric mode designates AI as the primary agent in the collaboration, where the AI system leads decision-making processes and operates with minimal human intervention. This mode often features automated interactions where the AI executes tasks independently, aimed at enhancing system capabilities and overall efficiency. Such a setup focuses on improving AI functionalities, maximizing system efficiencies, and generally involves a one-directional interaction from AI to human, primarily concentrating on tasks suited for AI’s computational abilities.
Recent research underscores the evolving role of AI in traditional human-machine settings. Liu and Zhao [33] discuss the allocation of roles in Human-Machine Collaboration (HMC), advocating for optimal strategies that seamlessly integrate intelligent systems into human-centric operations. Similarly, Wang et al. [34] examine the shift from human-human to human-AI collaboration, stressing the importance of mutual goal setting and collaborative task management to foster effective partnerships.
In this section, we looked at the fundamentals of Human-AI Collaboration, such as its main components and the various modes of operation. The evaluation of HAIC is critical to ensure that these systems are successful, efficient, and ethical.
The following part looks into the literature on HAIC evaluation, reviewing existing approaches and identifying gaps that suggest areas for further research.
2.3 Evaluation Approaches in HAIC
The evaluation of Human-AI Collaboration (HAIC) is essential to understanding the efficacy of these systems, identifying areas for improvement, and ultimately unlocking their full potential. This section examines current evaluation approaches in HAIC research. While Table 1 provides an overview of a diverse sample of studies across different domains, the analysis in this section will focus on selected key references to illustrate specific evaluation methodologies and their implications.
Through this analysis of a wide array of studies, our objectives are as follows:
  • Examine the advantages and drawbacks of current evaluation methodologies.
  • Identify the distinct difficulties presented by various HAIC applications and domains.
  • To effectively address the intricate and diverse aspects of HAIC in different sectors.
Table 1: Examples of Evaluation of Human-AI Collaboration Across Sectors
Reference Domain AI Technology/Method Qn Ql Mix Additional Info
Vollmuth et al. (2023) [35] Healthcare (Oncology) Image Analysis Accuracy is measured by sensitivity, precision, recall, F1-score. Focuses on human specialists to evaluate AI and HAIC performance.
Van Leeuwen et al. (2022) [36] Healthcare (Radiology) Image Analysis Emphasizes efficiency using sensitivity, specificity, and outcome metrics.
Timmons et al. (2023) [37] Healthcare (Mental Health) Not specified Investigates bias through interviews and case studies.
Farić et al. (2024) [38] Healthcare (Radiology) Diagnostic Support Systems Assesses early AI use via qualitative feedback from radiologists.
Calisto et al. (2022) [39] Healthcare (Breast Cancer) Medical Image Analysis Focuses on human-AI interaction through qualitative methods.
Rezwana et al. (2023) [40] Design Generative AI Develops interaction design framework using user feedback.
Tschandl et al. (2020) [8] Healthcare (Dermatology) Image Analysis Uses quantitative success metrics in high-stakes environments.
Siu et al. (2021) [41] Game-Based Reinforcement Learning Uses game performance metrics in Hanabi game as a testbed.
Reddy et al. (2021) [42] Healthcare Not specified Proposes a qualitative framework focusing on ethics and user perspectives.
Sharma et al. (2023) [43] Mental Health Reinforcement Learning Measures conversational empathy improvements through qualitative feedback.
Yik et al. (2024) [44] General (Benchmarking) Neuromorphic Computing Provides quantitative framework for unique AI context.
Rastogi et al. (2023) [45] General (Auditing) LLM Examines auditing collaboration using qualitative assessments.
Zhao et al. (2023) [46] Multimodal Interaction LLM with Visual Input Studies interaction through multimodal input performance metrics.
Makarius et al. (2020) [1] General (Business) Not specified Emphasizes social factors using qualitative methods.
Aleid et al. (2023) [47] Healthcare (Radiology) Image Analysis Focuses on accuracy using quantitative measures.
Al-Fatlawi et al. (2023) [48] Finance (Fraud) Genetic Algorithms Evaluates detection rates (true/false positives).
Yang et al. (2022) [49] General (Teaming) Behavior Modeling Uses quantitative and qualitative data to model human behavior for better interaction.
Aher et al. (2023) [50] Simulation-Based Large Language Models Simulates human behavior using quantitative metrics.
Verheijden & Funk [51] (2023) Design Generative AI Enhances design processes through qualitative co-creation feedback.
Hauptman et al. (2023) [52] General (Teaming) Adaptive AI Investigates adaptability perceptions through qualitative assessments.
Sankaran et al. (2022) [53] Manufacturing Recurrent Neural Networks Measures process performance using quantitative metrics.
Massaro (2022) [54] Manufacturing (Safety) Artificial Neural Networks Evaluates safety and risk through quantitative measures.
Oravec (2023) [55] Education Large Language Models Discusses academic integrity using qualitative methods.
El-Zanfaly et al. (2022) [56] Design/Art Physical Interface Focuses on creativity and interaction through qualitative feedback.
Arias-Rosales (2022) [57] Design/Art Shape Generation Combines quantitative shape metrics with qualitative value assessments.
Dziorny et al. (2022) [58] Healthcare (ICU) Decision Support Systems Examines high-stakes challenges using qualitative feedback.
Vössing et al. (2022) [59] General (Forecasting) LightGBM, Neural Networks Investigates AI transparency impact using mixed trust/performance metrics.
Chakravorti et al. (2022) [60] Finance Prediction Models Combines detection rates with qualitative trust assessments for AI markets.
Huang & Rust (2022) [61] Marketing Machine Learning Develops collaborative AI framework using mixed methods.
Sachan et al. (2024) [62] Finance Evidential Reasoning Reduces decision noise through quantitative measures.
Basu et al. (2021) [63] Finance (ICO) Linear Regression Studies decision-making using mixed quantitative and qualitative data.
Dikmen & Burns (2022) [64] Finance (Peer-to-Peer Lending) Explainable AI Examines trust and performance using mixed domain knowledge assessments.
Ma et al. (2024) [65] General (Decision-Making) Large Language Models Proposes deliberative AI using mixed decision-making metrics.
Alon-Barkat & Busuioc (2023) [66] Public Sector Regression Models Explores public sector interactions through mixed methods.
Nikolopoulou (2024) [67] Education Generative AI Assesses generative AI’s impact on pedagogy using mixed methods.
2.3.1 Advancements in Tools, Methodologies, and Real-Time Collaboration for HAIC
Human-AI Collaboration (HAIC) is increasingly evaluated using both traditional and innovative metrics, with an emphasis on ethical concerns, scalability, and adaptability across domains. Recent advancements in tools and methodologies are reshaping the evaluation landscape, offering novel ways to understand and optimize human-AI interactions.
Fischer (2023) [68] envisions the future of human-AI collaboration as a deeply intertwined partnership, where AI systems are designed not only to assist with tasks but also to foster positive emotional and social interactions. His work explores scenarios where AI colleagues are capable of understanding and responding to human emotions, aiming to enhance collaboration by creating a sense of trust and mutual satisfaction. This forward-looking perspective emphasizes the need for evaluation frameworks that address not only technical performance but also the social and emotional dimensions of human-AI interactions. Fischer’s work highlights the importance of incorporating human-centered design principles into AI systems to promote effective collaboration and user well-being.
The incorporation of ethical factors into evaluation frameworks is a growing trend, as evidenced by the work of Huang & Rust (2022) [61] in marketing and Nasir, Khan, & Bai (2023) [69] in healthcare. These frameworks emphasize the importance of transparency (e.g., disclosing the use of AI in decision-making), fairness (e.g., mitigating bias in algorithms), accountability (e.g., establishing clear lines of responsibility for AI-related outcomes), and human-centered design (e.g., prioritizing user needs and values) in AI implementation. Lase & Nkosi (2023) [70] further advocate for a Human-centric perspective in AI development, stressing the significance of aligning AI systems with user values and ethical principles. They propose using surveys and interviews to assess user perceptions of fairness, trust, and the impact of AI on their work.
Bojic et al. (2023) [71] highlight the need for standardized and robust evaluation practices, proposing a hierarchical framework to guide the design of reliable human evaluations of AI systems. This framework outlines a step-by-step process for defining evaluation objectives, selecting appropriate metrics (e.g., accuracy, efficiency, user satisfaction), designing experimental protocols, and analyzing data to draw valid conclusions about the effectiveness of HAIC.
Technological advancements further enable real-time monitoring and adaptive collaboration in HAIC. Recent studies, such as Woelfle et al. (2024) [11], emphasize the importance of benchmarking tools that evaluate human-AI collaboration through diverse performance metrics. These metrics extend beyond traditional KPIs like task completion time and error rates, incorporating dimensions such as evidence appraisal quality and the adaptability of AI systems to human inputs. Such tools provide actionable feedback on collaboration efficacy, fostering dynamic adaptability and enhancing the overall effectiveness of HAIC systems.
Interactive machine learning (iML) interfaces, explored by Saha et al. (2023) [72], represent a significant leap in fostering transparency and user engagement. By providing explanations for AI-generated recommendations and allowing users to adjust system parameters, iML interfaces enhance user comprehension, trust, and satisfaction. Saha et al. evaluate these interfaces using a combination of user surveys, interviews, and task-based performance metrics, demonstrating their potential to facilitate effective human-AI partnerships.
Together, these advancements underscore the importance of integrating ethical considerations, real-time analytics, and interactive tools to develop comprehensive evaluation frameworks for HAIC. These tools and methodologies not only enhance our understanding of human-AI collaboration but also pave the way for more adaptive, transparent, and ethical partnerships between humans and AI systems.
2.3.2 Analysis of Existing Approaches
In Table 1, we show a summary of Human-AI Collaboration (HAIC) evaluation practices we examined for this analysis. While quantitative methods remain dominant, there is a growing recognition of the value of qualitative and mixed-methods approaches for capturing the complex interplay between objective performance and subjective experiences.
Quantitative Focus: Objectivity and Measurable Outcomes
Quantitative evaluations, prioritizing objective measures to gauge system performance and efficacy,including the combination of humans and AI, are prevalent in HAIC research. In the medical field, Vollmuth et al. (2023) [35] and Van Leeuwen et al. (2022) [36] use tests like sensitivity, specificity, precision, recall, and F1-score to carefully check how well AI-assisted tools in oncology and radiology can make diagnoses. To compare how well AI can find fraud in the financial world, Al-Fatlawi et al. (2023) [48] use detection rates (true positives and false positives), and Sachan et al. (2024) [62] use quantitative measures to reduce decision noise in financial underwriting. Manufacturing studies, such as those by Sankaran et al. (2022) [53] and Massaro (2022) [54], also rely heavily on quantitative metrics to assess process performance and safety compliance, respectively. Researchers have developed quantitative frameworks to assess performance in diverse fields such as game-based AI (Siu et al., 2021) [41] and neuromorphic computing (Yik et al., 2024) [44] based on specific domain-relevant metrics.
While these quantitative approaches provide valuable insights into system performance and measurable outcomes, their focus on objective data often overlooks the nuanced human factors that influence the adoption and success of HAIC systems.
Qualitative Insights: Illuminating the Subjective Dimension
To complement the quantitative focus, a growing number of studies are embracing qualitative methods to delve into the subjective experiences of users and stakeholders. Timmons et al. (2023) [37] investigate biases in mental health AI through interviews and case studies, shedding light on potential ethical concerns and user perceptions. Their study reveals potential ethical concerns, such as the perpetuation of biases and the varying levels of trust users place in AI systems. Rezwana et al. (2023) [40] and Verheijden & Funk (2023) [40] use qualitative assessments in the design field to understand the impact of generative AI on creativity and design co-creation processes. Through focus group discussions and in-depth interviews with designers, these studies highlight how generative AI tools influence creative workflows. They discovered that while AI can enhance creative output by providing novel ideas, it also raises concerns about authorship and originality, which are critical to the design profession.
Sharma et al. (2023) [43] look at how conversational empathy improves in mental health applications using qualitative measures. Reddy et al. (2021) [42] suggest a qualitative framework to evaluate the use of AI in healthcare, with a focus on ethical issues and user perspectives. Their framework includes structured interviews with healthcare providers and patients, and thematic analysis to identify key concerns such as data privacy, trust in AI diagnoses, and the perceived impact on the patient-provider relationship. Their study underscores the need for AI systems to align with ethical standards and user expectations to be effectively integrated into healthcare practices.
By capturing the nuanced perspectives of users and stakeholders, these qualitative studies offer a rich understanding of the human factors that shape the adoption and impact of HAIC. However, the subjective nature of qualitative data can pose challenges in terms of generalizability and requires substantial resources for data collection and analysis.
2.3.3 The Mixed-Methods Promise: Towards a Holistic and Contextualized Evaluation
The limitations of relying solely on quantitative or qualitative approaches have led to the adoption of mixed-methods evaluations, which integrate objective performance metrics with subjective experiences. This approach provides the flexibility needed to tailor evaluations to specific domains while balancing the rigor of quantitative data with the depth of qualitative insights.
Domain-Specific Applications
In creative domains such as design and art, Arias-Rosales (2022) [57] combined quantitative shape generation metrics with qualitative assessments of designers’ perceived value. This dual approach highlighted that while AI-generated designs were technically innovative, their subjective aesthetic value varied significantly among designers, underlining the critical role of qualitative judgment in evaluating creative outputs. Similarly, in finance, Chakravorti et al. (2022) [60] evaluated AI prediction markets by pairing detection rates with trader and analyst perceptions. Their mixed-methods approach revealed that high prediction accuracy did not necessarily translate into trust, exposing concerns about transparency and interpretability.
Generalizable Frameworks
Mixed methods also demonstrate utility across domains. For instance, Yang et al. (2022) [49] modeled human behavior by combining task completion rates and error frequencies with user interviews, while Vössing et al. (2022) [73] integrated system accuracy metrics with trust assessments, uncovering the significant impact of transparency on user acceptance. In marketing, Huang & Rust (2022) [61] employed a mixed-methods framework that combined performance metrics, such as click-through rates, with ethical and user satisfaction insights. Together, these studies illustrate how mixed methods enable comprehensive evaluations that capture both technical performance and contextual user perceptions.
2.3.4 Tailored Evaluation for Domain-Specific Challenges
HAIC’s diverse application domains present unique challenges that demand tailored evaluation approaches, reflecting both technical complexities and the distinct priorities of stakeholders. As discussed in the Appendix A of this work, the evaluation of creative and linguistic AI requires tailored methodologies due to its inherent complexities.
Healthcare: Prioritizing Accuracy, Safety, and Ethical Considerations
In the healthcare, diagnostic accuracy and patient safety are critical. Studies like Vollmuth et al. (2023) [35] and Van Leeuwen et al. (2022) [36] use metrics like sensitivity, specificity but the high-stakes nature of medical decisions require additional qualitative approaches like interviews and focus groups (Farić et al., 2024 [74]; Calisto et al., 2022 [39]) to capture user perceptions, identify biases (Timmons et al., 2023 [37]), and ensure alignment with ethical principles (Reddy et al., 2021 [42]).
Creative Arts: Balancing Creativity, User Experience, and Ethical Impact
In creative arts, quantitative metrics (e.g. speed, output quality) are complemented byqualitative methods such as interviews and case studies (Rezwana et al., 2023 [40]; Verheijden & Funk, 2023 [51]) to asses AI’s influence on creativity and collaboration. Ethical concerns regarding bias and the potential undermining of human creativity, underscore the need for a balanced evaluation [75].
Finance: Mitigating Risks, Ensuring Trust, and Balancing Quantitative and Qualitative Insights
In finance quantitative metrics such as detection rates and false positives are crucial for fraud detection (Al-Fatlawi et al., 2023 [48]). However, understanding AI’s impact on decision-making and market stability also requires qualitative insights through surveys and interviews (Chakravorti et al., 2022 [60]). Mixed-methods approaches, provide a comprehensive view of human-AI interaction in financial markets. For a detailed discussion on the challenges specific to creative and linguistic AI, please refer to Appendix A.
In summary, while evaluation approaches vary significantly across domains, a comprehensive, mixed-methods framework is still needed to sully assess HAIC’s effectiveness and potential.
2.4 Towards a Unified Evaluation Framework
While existing HAIC evaluation approaches offer valuable insights, they remain fragmented. A unified framework is needed to integrate objective performance metrics with subjective experiences across diverse domains. Such a framework would support cross-study comparison, enhance generalizability, and standardize evaluation practices.
This framework should adapt to different industries, ensuring AI-assisted tools are both effective and aligned with human needs. In healthcare, it could integrate diagnostic accuracy with user trust; in creative fields, it could assess AI’s impact on artistic processes; and in finance, it could combine risk metrics with transparency evaluations. By addressing methodological gaps, this approach will help harness AI’s potential while ensuring its responsible integration into human workflows.
3 A Structured Framework for Human-AI Collaboration
Unlike previous studies that focus on isolated aspects of HAIC evaluation, the proposed framework provides a structured approach by setting measurable goals, clarifying tasks and reporting metrics [bolman2013reframing], that integrates multiple evaluation dimensions, organized in three primary factors and specialized subfactors. The primary factors are:
  • Goals – Evaluates individual and collective objectives within HAIC systems, with key metrics such as learning curves and prediction accuracy [76, 59].
  • Interaction – Assesses communication mechanisms, feedback processes, adaptability, and trust. Effective communication and transparency are critical for ensuring human-AI trust and usability [77].
  • Task Allocation – Examines role distribution, efficiency, and decision-making processes, ensuring that AI contributions complement human expertise in dynamic environments [78].
Each factor is further divided into measurable subfactors and associated metrics, allowing for a robust quantitative and qualitative assessment of HAIC effectiveness. To enhance practical applicability, we develop a decision-tree-based evaluation approach. This structured method assists practitioners in selecting relevant evaluation criteria based on the nature of the HAIC system. The decision tree ensures flexibility, allowing for tailored assessments in various domains, including healthcare, finance, manufacturing, and education.
By offering a holistic assessment tool, this framework addresses the key gaps identified in the literature and establishes a standardized methodology for evaluating HAIC in both research and industry applications.
3.1 Subfactors & Metrics of the Framework
Table 2 summarizes the detailed breakdown of factors to subfactors and their associated metrics.
Table 2: Factors and Subfactors of Human-AI Collaboration with Associated Metrics for Evaluation
Primary Factor Subfactor Human-AI Collaboration Metrics for Evaluation
Goals Individual Goals AI’s Goal: Learn Efficiently, Increase Accuracy. Human’s Goal: Provide Effective Teaching, Achieve Task Objectives. * Learning Curve (rate of improvement over time), * Prediction Accuracy (percentage of correct predictions), * Teaching Efficiency (time/resources spent per unit of learning) [76, 79, 80].
Collective Goals Shared Goal: Maximizing the overall performance of the AI system. * Overall System Accuracy (percentage of correct outputs), * Objective Fulfillment Rate (percentage of tasks successfully completed) [59, 81].
Interaction Communication Methods Query and Response Mechanism. * Ease of Use (user satisfaction with communication methods), * Clarity of Communication (understandability of queries and responses) [82, 83]
Feedback Mechanisms Human feedback on AI queries, AI adjustments based on feedback. * Frequency of Feedback (number of feedback instances), * Feedback Quality (relevance and specificity of feedback), * Feedback Impact (measured change in AI performance due to feedback) [83, 84, 82]
Adaptability AI’s ability to adapt query strategy and learning based on human input. * Strategy Adaptation Effectiveness (improvement in query selection based on feedback), * Adaptability Score (composite metric of flexibility and responsiveness) [82, 85, 86]
Trust and Safety Trust in AI system, ensuring user safety. * Trust Score (level of user trust in AI system) [87], * Safety Incidents (number of safety-related issues) [88], * Error Reduction Rate (decrease in errors due to AI assistance) [2], * Confidence (level of user trust in AI recommendations) [89]
Task Allocation Complementarity AI handles data analysis and uncertainty estimation; Humans provide expertise and clarification. * Query Efficiency (amount of information gained per query), * Expertise Utilization (degree to which human expertise is leveraged) [78]
Flexibility Adjusting query strategy based on interaction dynamics. * Adaptability Score (same as in Interaction), * Dynamic Task Allocation (ability to adjust task distribution based on real-time needs) [23]
Efficiency Streamlining the learning process through targeted queries. * Model Improvement Rate (rate of improvement in AI performance) [59, 78], * Learning Curve (same as in Goals), * Resource Utilization (efficiency in using computational resources) [59]
Responsiveness AI’s ability to rapidly adjust to new inputs and feedback. * Response Time (time taken to respond to queries or feedback), * Real-time Performance (ability to handle time-sensitive tasks) [82, 90]
Collaborative Decision Making Joint decisions on data labeling and correction of AI’s model. * Impact of Corrections (measured improvement in system performance through human intervention) [91, 92], * Decision Effectiveness (impact of decisions on overall task success) [93, 94]
Continuous Learning AI’s ability to learn from human input and adjust its model. * Model Improvement Rate (same as in Efficiency), * Learning Curve (same as in Goals), * Knowledge Retention (ability to retain learned information over time) [95]
Mutual Support AI’s ability to provide support to humans in the learning process; Humans support by providing information. * Error Reduction Rate (decrease in errors due to AI assistance) [2], * Confidence (level of user trust in AI recommendations) [96], * Task Completion Time (with and without AI support) [97]
Robustness Ensuring AI system’s robustness and reliability. * Adversarial Robustness (resilience to adversarial attacks) [98], * System Reliability (overall system uptime and fault tolerance) [99]
Goals Factor
The Goals factor ensures that the collaboration has a clear direction and that both human and AI efforts are aligned towards shared outcomes. This factor encompasses two key subfactors: Individual Goals, which pertain to the specific aims of each participant, and Collective Goals, which represent shared objectives that maximize the overall performance of the HAIC system.
Table 3 below summarizes the metrics used to evaluate these subfactors, providing concise definitions, measurement methods, and relevant references to support the evaluation of the Goals factor in Human-AI Collaboration (HAIC) systems.
Table 3: Summary of metrics for the Goals factor in Human-AI Collaboration.
Subfactor Metric Definition and Measurement
Individual Goals Learning Curve Measures the rate of improvement over time, indicating the effectiveness of human teaching and AI learning. This is plotted by graphing AI performance (e.g., accuracy) over iterations or training samples. A steep learning curve suggests effective learning [76].
Prediction Accuracy The proportion of correct predictions made by the AI model out of all predictions. This can be evaluated using metrics like accuracy, F1 score, or area under the ROC curve, depending on the task [80].
Teaching Efficiency The gain in AI performance per unit of teaching effort by the human. This is measured by tracking improvements in AI accuracy or error rate after human input, as explored in studies such as Cakmak and Thomaz [79].
Collective Goals Overall System Accuracy Calculated as a weighted average of performance metrics, reflecting the relative importance of each task in achieving the overall goal. This metric evaluates the general accuracy of the AI-human collaborative system, tailored to specific objectives [59].
Objective Fulfillment Rate The percentage of predefined goals successfully achieved within the collaboration. Often used in project management contexts to track progress and adapted to assess HAIC success [81].
Schneider et al. [93] provide empirical evidence on quality goal rankings, while Wang et al. [34] highlight the necessity of goal alignment in collaborative environments. Fiebrink et al. [100] illustrate how human input influences AI’s learning processes, emphasizing the mutual dependence of individual and collective goals.
Interaction Factor
The Interaction factor underscores the critical communication mechanisms through which humans and AI systems exchange information, provide mutual feedback, and adaptively respond to each other’s inputs. Key subfactors within this factor include Communication Methods, Feedback Mechanisms, Adaptability, and Trust and Safety. These subfactors capture the essential elements of interaction in Human-AI Collaboration (HAIC), ensuring seamless and effective communication, the incorporation of feedback, and adaptability to dynamic conditions.
Table 4 summarizes the metrics used to evaluate the Interaction factor, providing concise definitions and measurement methods for each subfactor.
Table 4: Summary of metrics for the Interaction factor in Human-AI Collaboration.
Subfactor Metric Definition and Measurement
Communication Methods Clarity of Communication Assesses how understandable the communication is between humans and AI. This can be evaluated using qualitative methods like expert review or user feedback and quantitative measures like the proportion of correctly interpreted queries or time taken to resolve ambiguities [101, 86].
Ease of Use Often measured through Likert scale questionnaires or user interviews to gauge the user’s subjective experience with the system.
Feedback Mechanisms Frequency of Feedback Monitors how often users provide feedback, helping developers assess engagement levels and identify areas for improvement [102].
Feedback Quality Measures the relevance and specificity of feedback, either qualitatively through expert reviews or quantitatively by analyzing the correlation between feedback and subsequent AI performance improvements. Hein et al. [103] explore feedback acceptance and motivation when provided by AI versus human supervisors.
Feedback Impact Quantifies the tangible effect of feedback on AI behavior or output quality, often measured by improvements in accuracy, precision, or other relevant metrics [104].
Adaptability Adaptability Score Combines measures of flexibility and responsiveness to assess how quickly and effectively the AI adapts to changes, as studied by El-Assady et al. [86].
Strategy Adaptation Effectiveness Measures improvement in AI’s query selection based on human feedback, incorporating insights from semantic interaction research [85].
Trust and Safety Trust Score Assesses user trust in the AI system through surveys and trust scales [105].
Safety Incidents Tracks the number of safety-related issues encountered to evaluate the reliability and safety of the AI system.
Ethical considerations such as bias, fairness, and transparency are critical to the success of HAIC. Significant research has underscored the complexity of achieving fair and unbiased AI systems. For example, Barocas et al. [106] discuss the ethical implications of bias in ML systems, highlighting the need for transparent and accountable AI practices. Holstein et al. [107] emphasize the importance of fairness in AI design and implementation, advocating for continuous monitoring and evaluation to mitigate bias.
Task Allocation Factor
Task allocation in HAIC focuses on strategically distributing responsibilities between humans and AI to leverage the strengths of each party and optimize overall performance. Effective task allocation enhances efficiency, minimizes errors, and increases user satisfaction, making it a critical component in the success of HAIC systems [108].
Table 5 summarizes the key subfactors and metrics associated with task allocation, providing concise definitions and measurement methods.
Table 5: Summary of metrics for the Task Allocation factor in Human-AI Collaboration.
Subfactor Metric Definition and Measurement
Complementarity Query Efficiency Indicates the number of queries the AI needs to reach a certain level of accuracy. Fewer queries signify more efficient learning [78].
Expertise Utilization The degree to which human expertise is leveraged. Measured qualitatively through observation or interviews, or quantitatively through metrics like the percentage of decisions where human input is decisive [19].
Flexibility Dynamic Task Allocation The ability to adjust task distribution based on real-time needs. Assessed by how quickly and effectively the system adapts to changes in workload, priorities, or resources [109].
Efficiency Model Improvement Rate The rate of improvement in AI performance. Calculated as the difference in performance between two time points, divided by the time interval. Applied to metrics like accuracy or error rates [59].
Resource Utilization The efficiency of computational resource use (e.g., CPU time, memory). Measured as a percentage of utilized resources compared to the total available.
Learning Curve The rate of improvement in AI performance over time, as previously defined in the Goals section.
Responsiveness Response Time The average time taken for AI or human participants to respond to queries or feedback. A lower response time indicates quicker and more effective communication [110, 82].
Real-time Performance The ability to handle time-sensitive tasks, measured by factors like throughput (tasks completed per unit time) and latency (delay in response) [90].
Learning Curve The rate of improvement in AI performance over time, as defined earlier.
Collaborative Decision Making Impact of Corrections Measures the improvement in system performance from human intervention. Quantified by changes in accuracy or error rates after corrections [92, 91].
Decision Effectiveness The impact of decisions on task success. Measured by comparing outcomes of decisions made with and without AI collaboration [94, 93].
Continuous Learning Knowledge Retention Assesses the AI’s ability to retain learned information over time. Evaluated by testing performance after periods of non-use [111].
Model Improvement Rate Same as defined in the Efficiency section.
Learning Curve Same as defined in the Goals section.
Mutual Support Error Reduction Rate Measures the decrease in errors made by the AI following human intervention. Demonstrates the effectiveness of human corrections [2].
Confidence The level of user trust in AI recommendations. Measured using Likert-scale questionnaires or surveys [105].
Task Completion Time Compares task completion times with and without AI support. Helps assess efficiency gains from HAIC [97].
Robustness Adversarial Robustness The resilience of the AI system to adversarial attacks. Measured through testing against adversarial examples.
System Reliability The overall system uptime and fault tolerance. Evaluated through system logs and reliability tests.
This structured approach, detailing subfactors and metrics, ensures a comprehensive evaluation of HAIC systems, enabling continuous improvement and effective collaboration between humans and AI.
The Table 6 offers definitions and calculations for each metric described in Table 2,facilitating a full knowledge of the assessment metrics used in the structured framework for Human-AI Collaboration (HAIC). These metrics are essential for quantitatively evaluating the efficacy, efficiency, and adaptability of HAIC systems across multiple domains.
Table 6: Evaluation Metrics for Human-AI Collaboration
Metric Formula/Description Example
Learning Curve Graphical representation of performance vs. number of training samples. In a medical image analysis task, a graph shows the AI’s accuracy in diagnosing tumors steadily increasing over 1000 training images.
[…]
Feedback Impact Impact = $\text{Performance Post-Feedback}-\text{Performance Pre-Feedback}Performance Post-Feedback - Performance Pre-Feedback$ A language translation AI’s BLEU score improves from 0.65 to 0.72 after incorporating user feedback on mistranslations.
Adaptability Score Adaptability Score = $\text{Performance Post-Adaptation}-\text{Performance Pre-Adaptation}Performance Post-Adaptation - Performance Pre-Adaptation$ A self-driving car adapts its driving style based on weather conditions, improving safety scores by 12%.
[…]
Impact of Corrections Improvement = $\text{Performance Post-Correction}-\text{Performance Pre-Correction}Performance Post-Correction - Performance Pre-Correction$ Human corrections to an AI writing assistant’s suggestions result in a 15% improvement in the text’s readability score.
[…]
Task Completion Time Time Saved = Time Without $AI\text{Time Without AI}-\text{Time With AI}Time With$out AI - Time With AI An AI-powered manufacturing robot reduces the time to complete a specific task from 10 minutes to 5 minutes.
[…]
Safety Incidents Safety Incidents = Number of safety-related issues encountered A self-driving car system logs 2 safety incidents per 1000 miles driven.
[…]
3.2 Building the Framework & Quantifying the Evaluation of HAIC
Building upon influential research by Fischer [112], Wenskovitch and North [85], and Sharma et. al [43], this framework offers a structured approach to Human-AI Collaboration (HAIC) evaluation. This is achieved using a decision tree (Fig. 2), that guides users through a series of questions to identify the most relevant metrics for their specific HAIC system. This combines the factors, subfactors, and metrics identified in the previous sections, based on the HAIC modes described earlier in Section 2.2, i.e., AI-Centric, Human-Centric, and Symbiotic.
Figure 2: Decision Tree Framework for Evaluating Human-AI Collaboration
Notice in Fig. 2 that each metric has a distinct color based on the main factor that is coming from the Table 2. This color-coding system helps users quickly identify the primary factor each metric belongs to, facilitating a more intuitive and user-friendly evaluation process. The framework is designed to be adaptable and scalable, accommodating a wide range of HAIC systems and applications. By providing a structured and comprehensive evaluation approach, this framework aims to enhance the understanding and assessment of HAIC, fostering the development of more effective and collaborative human-AI systems.
Each metric is color-coded based on its primary factor (Table 2), enhancing clarity and usability. This adaptable, scalable framework fosters a comprehensive understanding of HAIC systems, improving their assessment and development.
AI-Centric Collaboration
Focuses on AI autonomy, emphasizing efficiency and uncertainty handling metrics.
Human-Centric Collaboration
Evaluates AI’s communication and support for human tasks, emphasizing usability and user satisfaction.
Symbiotic Collaboration
Assesses adaptability and task distribution to optimize human-AI strengths.
By structuring evaluation through this decision-tree approach, we enhance consistency, enabling better-informed HAIC system development and optimization. This approach simplifies evaluation, tailors assessments to HAIC systems, and remains adaptable as technologies evolve.
Weighting Mechanism & Overall Score Calculation
While the decision tree and associated metrics lead the evaluation process, we recognize the importance of a weighting mechanism to provide a more comprehensive evaluation of HAIC effectiveness. A weighting mechanism ensures a comprehensive evaluation, adjusting metric importance based on HAIC mode. Weighted averages determine overall scores, enhancing accuracy and adaptability. Prior research supports this approach, emphasizing context-specific evaluation [113][83][59]. Future work should refine these weights through empirical studies to strengthen HAIC assessment robustness.
4 Implications of the HAIC Evaluation Framework to Real-World Domains
Although it has not yet been empirically validated, the proposed HAIC framework offers theoritical basis for assesing human-AI collaboration across sectors such as manufacturing, healthcare, finance, and education. By emphasizing factors like Adaptability, Confindence and Clarity of Communication this framework lays the groundwork for future empirical studies.
4.1 Manufacturing
Manufacturing’s focus on safety, accuracy, and productivity aligns with the HAIC Symbiotic mode, which integrates human expertise with AI skills. Studies [114, 16, 115, 116] demonstrate a shift towards worker-centric, adaptive approaches that highlight critical metrics such as Adaptability Score, Error Reduction Rate, Confidence, and Task Completion Time.
4.2 Healthcare
In medical imaging diagnosis, AI supports human decision-making by enhancing diagnostic accuracy and efficiency [117, 118, 119, 47, 120]. The framework targets improvements in diagnostic Accuracy and Efficiency, emphasizing clear AI outputs. Evaluation parameters include System Accuracy, Prediction Accuracy, Response Time, and Decision Effectiveness, with further research needed on workflow impacts and patient communication.
4.3 Finance
The finance sector benefits from a Symbiotic approach where AI’s analytical capabilities complement human expertise to enhance fraud detection and operational efficiency [121, 48, 122, 123]. Key aspects include achieving goals of Accuracy and Efficiency and ensuring clear communication of AI insights. Metrics such as the Error Reduction Rate help assess system effectiveness, with future work needed to address ethical and long-term impacts.
4.4 Education
In education, a mix of Symbiotic and Human-Centric models enhances teaching and learning [124, 125, 34, 43, 10, 126]. The focus is on improving Teaching Efficiency and achieving personalized learning through effective communication. Important metrics include Task Completion Time and Learning Curve, while further research should explore fostering active student engagement and critical thinking.
5 Limitations in Investigating Behavioral Factors
Our framework for evaluating Human-AI Collaboration (HAIC) emphasizes goals, interaction dynamics, and task allocation while excluding behavioral aspects due to methodological and scope constraints. Investigating behavioral factors involves interdisciplinary approach from psychology, sociology, cognitive science, which require complex and varried methodologies. While these factors are crucial for understanding trust and collaborative goals over time—exemplified by studies such as Yang, Carroll, & Dragan (2022) [49]—our study focuses on structural elements that are more readily quantifiable. We advocate for future research to explore behavioral dimensions in-depth to complement our objective framework.
6 Conclusion
This paper makes a two-fold contribution to HAIC evaluation. First, it offers a comprehensive literature review of the existing methodologies and their limitations. Second, it introduces a structured framework that integrates key evaluation factors identified in the literature and provides a standardized method for assessing HAIC across various domains.
A key gap in current methodologies is the absence of a framework that integrates both qualitative and quantitative metrics across diverse sectors. The framework proposed herein addresses this limitation by offering a structured evaluation approach based on the mode of collaboration rather than a specific sector or AI application. By breaking down evaluation into smaller, manageable components, it provides a flexible method to assess HAIC across different contexts.
Although the framework is theoretical at this stage, future research will focus on empirical validation through software implementation and real-world applications across various sectors, ensuring its adaptability and robustness. Additionally, our study acknowledges limitations, particularly the exclusion of behavioral and ethical considerations from the evaluation framework. These dimensions, though outside the current scope, are essential aspects of HAIC that should be explored in future research.
Acknowledgments
The work leading to these results has received funding from the EU’s Horizon Europe research and innovation programme under GA No. 101120218, project HumAIne. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A The Unique Challenge of Evaluating Creative and Linguistic AI
Establishing comprehensive frameworks for assessing HAIC is vital, however certain applications pose distinct challenges due to their unique nature. Two prominent examples are Large Language Models (LLMs) and Generative AI in the Arts, where AI’s impact on human creativity, expression, and communication requires specialized evaluation methodologies.
While our framework provides a foundation, traditional metrics may be inadequate for these domains. Can quantitative measures fully capture artistic collaboration or AI’s influence on human language? How do we assess ethical concerns, bias, or AI-generated content’s societal impact? These challenges call for a nuanced evaluation approach. Although we don’t address these challenges in this work, we briefly discuss them here for completeness.
A.1 Large Language Models (LLMs)
LLMs have transformed text generation, translation, and dialogue. The GPT-3 [127] family, including ChatGPT and GPT-4, demonstrates their capacity for complex NLP tasks [128]. Evaluating HAIC with LLMs requires developing methodologies beyond traditional metrics [129].
Ensuring interpretability and fairness is vital. Binz & Schulz (2022)[130] show that even high-performing LLMs may use opaque reasoning patterns. Evaluation frameworks must assess how well LLMs explain their outputs and mitigate biases. Aher et al. (2022)[50] highlight the importance of testing LLM adaptability in user interactions.
LLMs augment cognition, enhance creativity, and personalize learning. Webb et al. (2022)[131] show their potential in analogical reasoning. Douglas (2023)[132] emphasizes their impact on ideation and artistic expression, while Liu et al. (2023) [133] explore their role in education. Evaluating these aspects requires metrics for creativity, knowledge retention, and learner autonomy.
Longitudinal studies provide insights into HAIC with LLMs. Yin et al. (2023)[134] suggest tracking how AI capabilities and user trust evolve over time. Dwivedi et al. (2023) [135] address societal implications, underscoring the need for ethical AI use.
Developing robust LLM evaluation frameworks is essential. Addressing interpretability, bias, and interaction quality will help create responsible, effective HAIC systems.
A.2 Generative AI in the Arts
Generative AI is reshaping artistic expression across visual art, music, and poetry. AI moves beyond tools to co-creators, expanding artistic possibilities. Notable examples include AI-generated paintings, compositions, and interactive installations. Gmeiner et al. (2022) [136] suggest leveraging team learning frameworks to enhance AI collaboration in creative fields.
Evaluating HAIC in generative AI requires adapting traditional art metrics—such as aesthetic quality, originality, and audience reception—to consider AI’s role in the creative process [137]. Verheijden & Funk (2023) [51] stress assessing AI’s impact on ideation, communication, and artistic workflows.
New methodologies must evaluate AI-assisted artistic creation holistically. He et al. (2023)[138] propose co-creativity as a joint search process, offering insights into human-AI dynamics. Ethical concerns also arise, including copyright, ownership, and authenticity. Epstein et al. (2023)[139] call for interdisciplinary research into AI’s cultural and economic impact.
As AI-assisted art evolves, evaluation models must adapt. Future research should examine real-time, personalized experiences and how AI reshapes social interactions in creative settings [140]. Documenting the artist-AI collaboration process may become as important as assessing final works.